I am suppose to solve the systems of equations by elimination and I don't understand it. Here is my problem.
X+y=-6
2x-y=3

Answer by mananth(16946)

1 x + 1 y = -6 .............1

2 x -1 y = 3 .............2
Eliminate y
multiply (1)by 1
Multiply (2) by 1
1 x 1 y = -6
2 x -1 y = 3
Add the two equations
3 x = -3
/ 3
x = -1
plug value of x in (1)
1 x + 1 y = -6
-1 + y = -6
y = -6 + 1
y = -5